OpenStudy (anonymous):

Estimate 8 and one-third times 2 and four-fifths using any method you like. Then find the actual product. Use your estimate to show that your answer is reasonable.

OpenStudy (anonymous):

First, change the fractions into improper fractions. \[\huge 8\frac { 1 }{ 3 } =\frac { 25 }{ 3 } \] \[\huge \frac { 24 }{ 5 } =\frac { 14 }{ 5 } \] \[\huge Numerator: 25 * 14 = 350\] \[\huge Denominator: 3 * 5 = 15\] ANSWER: 350/15 or 23.333 or 23 1/3

OpenStudy (anonymous):

Thank you.